MOLECULAR ORBITAL STUDY OF THE DISSOCIATIVE RECOMBINATION; HC_{3}NH^{+} + e^{−}. IS IT POSSIBLE TO PRODUCE ALL OF ISOMERS OF CYANOACETYLENE?

\underline{K. FUKUZAWA} AND Y. OSAMURA, Department of Chemistry, Faculty of Science, Rikkyo University, 3-34-1 Nishi-ikebukuro, Toshima-ku, Tokyo 171, Japan; H. F. SCHAEFER III, Center for Computational Quantum Chemistry, University of Georgia, Athens, GA 30602, USA.

The dissociative recombination reaction between HC_{3}NH^{+} and an electron is one of the major route to produce cyanoacetylene, HC_{3}N, in interstellar clouds. We have studied various pathways of this recombination reaction producing HC_{3}N and its isomers, HNC_{3}, HCCNC, and HCNCC, theoretically. Potential energy surfaces for the processes from neutralized HC_{3}NH are examined by using the ab initio molecular orbital method. The calculated result shows that HCCNC is also produced via the isomerization processes in addition to the products HC_{3}N and HNCCC from the direct hydrogen dissociation. Several product channels C_{2}H + HNC, C_{2}H + HCN, NH + C_{3}H, CN + C_{2}H_{2}, C_{3} + NH_{2} and C_{3}N + H_{2} are shown to be energetically possible based on the thermochemical relationships.